OpenStudy (anonymous):

Help me with L'hopital rule. lt x--->3 [(1/(x-3)) - (1/log(x-2))]

OpenStudy (anonymous):

@mathmate plz help @Psymon plz help its urgent

OpenStudy (psymon):

Well, the first step is to actually shift the function where it is in an undeterminant form, 0/0, infinity/infinity, etc.

OpenStudy (anonymous):

its 0/0....What is the differentiation of log x?

OpenStudy (psymon):

Ah, it looks like it's just two separate terms, I apologize. Well, the derivative of log(base a)u is [1/[lna(u)] * du/dx

OpenStudy (psymon):

I suppose I should rewrite that better. [1/[(lna)*u]] * du/dx

OpenStudy (mathmate):

To put it in 0/0 form, you can add the two fraction in the usual way.

OpenStudy (dumbcow):

\[\frac{1}{x-3} - \frac{1}{\ln (x-2)} = \frac{\ln(x-2)-(x-3)}{\ln(x-2) (x-3)}\] now differentiate \[\large \frac{\frac{1}{x-2} - 1}{\ln(x-2) + \frac{x-3}{x-2}}\]

OpenStudy (mathmate):

I get \( \frac{(log(x-2)-x+3)}{((x-3)log(x-2))} \) Now you can apply l'Hôpital rule (twice) to get your answer

OpenStudy (anonymous):

thanks a lot @mathmate @Psymon @dumbcow